# Learning phases with Quantum Monte Carlo simulation cell

### v1

@@ Abstract
We propose a new machine learning input data type, “spin-opstring”, derived from Stochastic Series Expansion Quantum Monte Carlo (QMC) simulations. It offers a compact, memory-efficient representation of QMC simulation cells, combining the initial state with an operator string that encodes the state’s evolution through imaginary time. Using supervised machine learning on two models, we demonstrate the input’s effectiveness in capturing both conventional and topological phase transitions. Additionally, we conduct a regression task to predict superfluid density, which reflects non-local properties of the quantum system, and achieve good accuracy. These results validate the spin-opstring as an effective input for machine learning applications.
@@ I Introduction
In the past decade, due to their ability to analyze complex, high-dimensional data and uncover patterns that are challenging to detect using traditional methods, neural networks and machine learning (ML) techniques have become an integral part of condensed matter physics research (for a review, see for e.g. [1, 2, 3]). In particular, classifying phases and learning information about them, using supervised and unsupervised methods, have been a central focus of the ML approach [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In this category of ML, various kinds of input data, such as spin configuration, correlation functions, entanglement spectra, winding numbers etc. are used to train the model [10, 16, 7, 11, 9, 17, 18]. In line with the philosophy of artificial intelligence, it is always better to apply as little preprocessing as possible to the training data such that the ML model is able to autonomously identify the relevant features.
There have been proposals to use the raw $(d+1)$-dimensional configuration space generated by quantum Monte Carlo (QMC) methods to train ML models for a $d$-dimensional system, where the additional dimension corresponds to the imaginary time $\beta=1/T$ (with $T$ being the temperature). For increasing system sizes and low temperatures, which is necessary to study quantum phase transitions, the imaginary time direction increases rapidly leading to large simulation cells, which is too large to be effectively handled by machine learning algorithms. In Ref. [10, 19], the auxiliary field configurations obtained from determinant QMC simulations was utilized successfully as ML input, however, for only small systems at high temperatures. To address the challenge of large simulation cells, the authors of Ref. [20] proposed a compression scheme for the $(d+1)$-dimensional space-time configurations generated by Stochastic Series Expansion QMC, to reduce their size to a manageable level and successfully used them as input for a convolutional neural network (CNN).
@@
During an SSE QMC simulation, the full set of propagated states is not stored on the computer memory. Instead, only one propagated state and the operator string are stored for each QMC update, from which the remaining states can be generated on the fly. This approach substantially reduces memory requirements compared to other world-line QMC methods [24], where entire space-time configurations are stored. Leveraging this unique feature of SSE, we propose using the initial state along with the operator string, which we call “spin-opstring”, as input data for ML models. In comparison to previous studies that have used QMC simulation cells as input, our input preserves all the information contained in a simulation cell while being memory-efficient.
@@
To demonstrate the effectiveness of the spin-opstring ML input, we applied supervised machine learning techniques to two different models. The first model is the bosonic $t_{2}-V_{1}$ model, which, as reported in Ref. [25], is the simplest model to realize a checkerboard supersolid phase on a two-dimensional square lattice. Using supervised ML with labeled spin-opstring input data generated by SSE QMC, we successfully reproduced the phase transitions of this model. Additionally, we performed a regression task with the ML model to predict the superfluid density, which is in close agreement to the QMC results. To further evaluate the performance of the spin-opstring data in the context of topological phase transitions, we applied supervised ML to the one-dimensional Su-Schrieffer-Heeger (SSH) model [26]. We found that despite the absence of an obvious order parameter, the spin-opstring data is able to capture the transition correctly and predicts the transition point accurately.
@@
The paper is organized in the following manner. In Sec. II we give a brief description of the SSE QMC algorithm and describe the spin-opstring input data. In Sec. III, we outline the input data format for machine learning, followed by a discussion on the neural network architecture and its training procedure. Next, we present the results obtained from the ML model in Sec. IV, and conclude with a summary of our findings in Sec. V.
@@ II Proposal
In the above expression, the summation over $n$ varies from 00 to $\infty$, resulting in operator strings or “opstrings” $S_{n}$ of varying length, which are numerically inconvenient to handle. To overcome this difficulty the Taylor expansion is truncated at some finite cutoff $M$, the value of which is dynamically chosen in the algorithm in such a way that the exponentially small truncation errors can be neglected. For the convenience in computation, $(M-n)$ identity operators, $H_{0,0}=I$, are added to a string of $n$ non-identity operators so that the length of all opstrings become equal. Now, $(M-n)$ identity operators can be inserted among the $M$ total operators in $\binom{M}{M-n}=\frac{M!}{(M-n)!n!}$ ways. To avoid over-counting due to the identical nature of the identity operators, we can divide the partition function by $\binom{M}{M-n}$, leading to the expression,
@@
The update procedures in SSE QMC algorithm ensure sampling of states with non-zero contributions to the trace while preserving ergodicity. It is achieved by employing two types of updates diagonal update and off-diagonal update. The diagonal update adjusts the number of non-identity Hamiltonian operators $n$ in the operator string $S_{M}$, by replacing diagonal operators with identity ones and vice-versa. On the other hand, off-diagonal update replaces certain diagonal operators with off-diagonal ones and vice-versa, without altering the total number of operators $n$ in the operator strings. For a pedagogical introduction to the various SSE update procedures including recent developments, see [27].
@@
Taking advantage of this specific feature of SSE, in this paper, we propose the initial state $|\alpha(0)\rangle$ along with the operator string $S_{M}$ as the input data for ML models. In contrast to previous studies, where simulation cells has been used as input of ML models, our input retains the entire information stored in a simulation cell, without being memory-extensive.
In SSE simulation, the $\uparrow$ and $\downarrow$ spins are usually stored as $+1$ or $-1$, represented by $\sigma(i)$ in Fig. 1, where $i$ indicates the site number. Moreover, in Fig. 1 the open and solid bars indicate diagonal and off-diagonal operators, respectively, while the absence of a bar between states correspond to a fill-in identity operator. The bond number on which an operator acts at level $p$ is denoted by $b(p)$. If a diagonal operator is present at level $p$, a value $1$ is assigned to the variable $a(p)$, whereas a value $2$ is assigned for an off-diagonal operator. Using this convention, an operator string $S_{M}$ can be constructed such that the diagonal and off-diagonal operators are encoded as even and odd integers, respectively, following the rule $S(p)=2b(p)+a(p)-1$, whereas for identity operators $S(p)$ is assigned to a value 00. We append this operator string $S_{M}$ after the initial spin configuration and prepare the one-dimensional “spin-opstring” input data.
For a system with $N_{s}$ number of sites and an operator string of length M, the storage requirements for the total simulation cell is $N_{s}(1+M)$ memory units, whereas for the spin-opstring data it is just $(N_{s}+M)$. As an example for a $12\times 12$ square lattice ($N_{s}=144$) with a typical $M=25000$, the spin-opstring requires about $140$ times less memory than the whole simulation cell.
@@ III Method
Before feeding the spin-opstring data to the machine, we preprocess the data by applying the following three steps. Firstly, the initial state $|\alpha\rangle$ is mapped from the spin basis to the occupation basis, such that $\uparrow$ spin at site $i$ ($\sigma(i)=1$) maps to the number operator $n_{i}=1$ and $\downarrow$ spin ($\sigma(i)=-1$) maps to $n_{i}=0$. Secondly, the operator string is normalized between values 00 and $1$, by dividing the operator string elements by the maximum possible value, which is $12N_{b}+$ 1, where $N_{b}$ is the total number of bonds in the system. Lastly, since the length of the operator string varies is SSE QMC simulations at different regions of the phase diagram, we determine the maximum length of the training as well as prediction data and apply a padding of zeroes at the end of the operator strings as required to equalize the length of the data. As a zero in the operator string signifies the identity operator, padding with zeroes at the end does not change the physical properties of the system encoded in the data.
@@
To analyze the one-dimensional spin-opstring data, we have used the usual feed-forward neural network architecture with fully-connected layers and/or one-dimensional convolutional neural network (CNN) layers. Each of these layers consists of an initial fully connected or convolution step, followed by batch normalization and a rectified linear unit (ReLU) activation function. For the final layer, a fully connected layer is used with a softmax activation function for classification tasks, and a ReLU activation for regression tasks, without batch normalization in both cases. During training, we have minimized the categorical cross-entropy loss function for classification tasks and the mean squared error for regression. The stochastic gradient descent learning process has been optimized by using the Adam algorithm [28]. The code has been implemented in Python by using the Keras library [29]. The ML models used in our study are schematically represented in Table 1 and are further described in detail below.
For the phase classification in the two-dimensional $t_{2}-V_{1}$ model, we have used a ML model with three fully-connected layers: the first input layer consisting of a number of neurons equal to the spin-opstring data length, followed by a hidden layer of 32 neurons, and finally an output layer consisting of a few neurons, depending on the number of phases, with softmax activation. In contrast, for regression, the ML model consists of two fully-connected hidden layers, each with $64$ neurons and the final layer has a single neuron with a ReLU activation function.
Classification model Regression model
FC $n=\mathrm{input}$ BatchNorm ReLU Conv1D $f_{n}=16$, $f_{s}=9$ BatchNorm ReLU FC $n=\mathrm{input}$ BatchNorm ReLU
@@
For phase boundary detection in the 1D SSH model, we have used a CNN architecture. After the input layer, this architecture includes $4$ hidden CNN layers with $16,32,64$ and $128$ filters of sizes $9,5,5$ and $5$, respectively. Each CNN layer is followed by a MaxPool layer with a stride of $2$, which reduces the length of the 1D data by half. The output of the last CNN layer is then passed to a fully-connected layer with two neurons, corresponding to the two phases of the SSH model. For each of the models discussed above, we present our results obtained as an average of the predictions from 100 independently trained instances of the same model.
IV Applications
@@ IV.1 Two-dimensional $t_{2}-V_{1}$ model
We first apply our proposal to the two-dimensional (2D) $t_{2}-V_{1}$ model which belongs to the family of extended Bose-Hubbard models in the hardcore limit. This simple model displays a rich phase diagram and is, in fact, the minimum model to realize lattice supersolidity of the checkerboard-type. Interestingly, the system also exhibits a line of mixed-order phase transitions with varying particle density which displays the extreme Thouless effect at half-filling. In the 2D $t_{2}-V_{1}$ model, a system of hard-core bosons (HCBs) is studied on a periodic square lattice described by the Hamiltonian,
@@
[…]
@@
Figure 7: ML prediction of the SSH model performed on lattice with $32$ sites. As $v$ is varied from 00 to $1$, $w$ changes according to $1-v$. The dashed line represents the phase boundary predicted by the ML model.

### v2

# thanks: Corresponding author

@@ Abstract
We propose the use of the “spin-opstring”, derived from Stochastic Series Expansion Quantum Monte Carlo (QMC) simulations as machine learning input data. It offers a compact, memory-efficient representation of QMC simulation cells, combining the initial state with an operator string that encodes the state’s evolution through imaginary time. Using supervised machine learning, we demonstrate the input’s effectiveness in capturing both conventional and topological phase transitions. Additionally, we conduct a regression task to predict superfluid density, which reflects non-local properties of the quantum system, and achieve good accuracy. We also demonstrate the capability of spin-opstring data in transfer learning by training models on one quantum system and successfully predicting on another, as well as showing that models trained on smaller system sizes generalize well to larger ones. Finally, using two state-of-the-art interpretability techniques, Layer-wise Relevance Propagation and SHapley Additive exPlanations, we show that the machine learning models learn and rely on physically meaningful features from the input data. Together, these findings establish the spin-opstring as a broadly-applicable and interpretable input format for machine learning in quantum many-body physics.
footnotetext: These authors contributed equally to this work.
@@ I Introduction
In the past decade, due to their ability to analyze complex, high-dimensional data and uncover patterns that are challenging to detect using traditional methods, neural networks and machine learning (ML) techniques have become an integral part of condensed matter physics research (for a review, see for e.g. [1, 2, 3]). In particular, classifying phases and learning information about them, using supervised [4, 5] and unsupervised [6, 7, 8, 9] methods, have been a central focus of the ML approach. More recently, the community has increasingly sought to interpret the decisions of ML models, in order to overcome their inherent “black box” nature (see [10] for a review). While various kinds of input data, such as eigenstates [11], correlation functions [12, 5], entanglement spectra [13, 8], winding numbers [9] etc. are used to train these models, their choice often require existing knowledge about the properties of the system along with analytical or numerical computations. In line with the philosophy of artificial intelligence, ML models should ideally autonomously identify relevant features with minimal data preprocessing.
Raw spin configurations obtained as snapshots from Monte Carlo simulations have been successfully used as ML input in phase classification and related tasks [4, 6]. However, they do not fully encode the quantum nature of the system which might be crucial for characterization of quantum phase transitions [14]. There have been proposals to use the raw $(d+1)$-dimensional configuration space generated by quantum Monte Carlo (QMC) methods to train ML models for a $d$-dimensional system, where the additional dimension corresponds to the imaginary time $\beta=1/T$ (with $T$ being the temperature). For increasing system sizes and low temperatures, which is necessary to study quantum phase transitions, the imaginary time direction increases rapidly leading to large simulation cells (usually $\mathcal{O}(10^{2})$ Gigabytes and higher), which is too large to be effectively handled by machine learning algorithms. In Ref. [15, 16], the auxiliary field configurations obtained from determinant QMC simulations were utilized successfully as ML input, however, for only small systems at high temperatures. To address the challenge of large simulation cells, the authors of Ref. [17] proposed a lossy compression scheme for the $(d+1)$-dimensional space-time configurations generated by Stochastic Series Expansion QMC, to reduce their size to a manageable level and successfully used them as input for a convolutional neural network (CNN). However, the above-mentioned QMC algorithm provides an in-built lossless encoding scheme which we shall utilize in this work.
@@
During an SSE QMC simulation, the full set of propagated states in a single update is not stored on the computer memory. Instead, only one propagated state and the operator string are stored for each QMC update, from which the remaining states can be generated on the fly. This approach substantially reduces memory requirements compared to other world-line QMC methods [21], where entire space-time configurations are stored. It would therefore be interesting to ask whether this unique compact representation of the simulation cell can act as a meaningful input for ML tasks. In this work, we propose to use as ML input the initial state (or first configuration) of the simulation cell combined with the operator string from each QMC update, which we call in short the “spin-opstring”. In comparison to previous studies that have used QMC simulation cells as input, our choice of input preserves all the information contained in a simulation cell while being memory-efficient.
@@
To demonstrate the effectiveness of the spin-opstring ML input for phase classification and regression tasks, we applied supervised machine learning techniques to two different models. The first model is the bosonic $t_{2}-V_{1}$ model, which, as reported in Ref. [22], is the simplest model to realize a checkerboard supersolid phase on a two-dimensional square lattice. Using supervised ML with labeled spin-opstring input data generated by SSE QMC, we successfully reproduced the phase transitions of this model. Additionally, we performed a regression task with the ML model to predict the superfluid density, which is in close agreement to the QMC results. To further evaluate the performance of the spin-opstring data in the context of topological phase transitions, we applied supervised ML to the one-dimensional Su-Schrieffer-Heeger (SSH) model [23]. We found that despite the absence of an obvious order parameter, the spin-opstring data is able to capture the transition correctly and predicts the transition point accurately.
We further explored the application of spin-opstring data for transfer learning tasks. First, we applied our ML model trained on data from the $t_{2}-V_{1}$ model to successfully predict a phase transition in the $t_{1}-V_{1}$ model. Next, we were able to predict phase transitions in the interacting 1D SSH model using a ML model trained in the non-interacting case. Additionally, we showed that it is possible to generalize a ML model trained with spin-opstring data from a smaller system to larger systems. To better understand the benefit of using the full spin-opstring instead of just the spin-state or operator string alone, we discuss some examples, where the full input proves indispensable.
Having established the usefulness of this input datatype, we turn our focus to the interpretation of the decision-making process in the trained ML models. It is important to mention here that the majority of studies dealing with interpretability put their emphasis on spin states as the input data [24, 25, 26]. The suitability of alternative input representations in such analyses remain mostly unexplored [27, 28]. To the best of our knowledge, our study marks the first instance of interpreting ML models trained on simulation cell data. To this end, we utilize two well-known interpretability methods, Layer-wise Relevance Propagation (LRP) [29] and SHapley Additive exPlanations (SHAP) [30] to understand which features in the data are most important to the model’s predictions. We developed a strategy to extract meaningful insights, based on different physical operators, from the large-dimensional feature scores obtained directly from LRP and SHAP. In the classification, regression and transfer learning tasks, we find that ML models trained with spin-opstring data are able to identify key physical features from each phase, showcasing their ability to learn and utilize essential quantum characteristics.
@@
The paper is organized in the following manner. In Sec. II we give a brief description of the SSE QMC algorithm and describe the spin-opstring input data. In Sec. III, we outline the input data format for machine learning, followed by a discussion on the neural network architecture and its training procedure. Next, we present the results obtained from the ML model in the context of phase classification and regression in Sec. IV. Sec. V explores further applications of the spin-opstring input in transfer learning. We compare the performance of using only the spin-state or the operator string with that of the combined spin-opstring in Sec. VI. Sec. VII focuses into understanding the model’s decision-making process through the use of established interpretability methods. Finally, we conclude with a summary of our findings in Sec. VIII.
@@ II Proposal
In the above expression, the summation over $n$ varies from 0 to $\infty$, resulting in operator strings or “opstrings” $S_{n}$ of varying length, which are numerically inconvenient to handle. To overcome this difficulty the Taylor expansion is truncated at some finite cutoff $M$, the value of which is dynamically chosen in the algorithm in such a way that the exponentially small truncation errors can be neglected. For the convenience in computation, (Mn)(M-n) identity operators, $H_{0,0}=I$, are added to a string of $n$ non-identity operators so that the length of all opstrings become equal. Now, (Mn)(M-n) identity operators can be inserted among the $M$ total operators in $\binom{M}{M-n}=\frac{M!}{(M-n)!n!}$ ways. To avoid over-counting due to the identical nature of the identity operators, we can divide the partition function by $\binom{M}{M-n}$, leading to the expression,
@@
The update procedures in SSE QMC algorithm ensure sampling of states with non-zero contributions to the trace while preserving ergodicity. It is achieved by employing two types of updates, diagonal update and off-diagonal update. The diagonal update adjusts the number of non-identity Hamiltonian operators $n$ in the operator string $S_{M}$, by replacing diagonal operators with identity ones and vice-versa. On the other hand, off-diagonal update replaces certain diagonal operators with off-diagonal ones and vice-versa, without altering the total number of operators $n$ in the operator strings. For a pedagogical introduction to the various SSE update procedures including recent developments, see [31]. It is important to note that, while SSE is fundamentally a finite-temperature QMC method, it can effectively capture a system’s ground state when a sufficiently large inverse temperature $\beta$ is chosen [32]. For the systems examined in this study, we have observed that setting $\beta$ within the range of $1.5L$ to $2L$, with $L$ representing the linear dimension of the system, reliably led to the system converging to its ground state.
@@
Taking advantage of this specific feature of SSE, in this paper, we propose to use the initial state $|\alpha(0)\rangle$ of the simulation cell in a given update, along with the operator string $S_{M}$ as the input data for ML models. In contrast to previous studies, where simulation cells have been used as input of ML models, our choice of input retains the entire information stored in a simulation cell, without being memory-extensive.
In SSE simulation, the $\uparrow$ and $\downarrow$ spins are usually stored as +1+1 or 1-1, represented by $\sigma(i)$ in Fig. 1, where $i$ indicates the site number. Moreover, in Fig. 1 the open and solid bars indicate diagonal and off-diagonal operators, respectively, while the absence of a bar between states correspond to a fill-in identity operator. The bond number on which an operator acts at level $p$ is denoted by $b(p)$. If a diagonal operator is present at level $p$, a value $1$ is assigned to the variable $a(p)$, whereas a value $2$ is assigned for an off-diagonal operator. Using this convention, an operator string $S_{M}$ can be constructed such that the diagonal and off-diagonal operators are encoded as even and odd integers, respectively, following the rule S(p)=2b(p)+a(p)1S(p)=2b(p)+a(p)-1, whereas for identity operators $S(p)$ is assigned to a value 0. We append this operator string $S_{M}$ after the initial spin configuration and prepare the one-dimensional (1D) “spin-opstring” input data.
For a system with $N_{s}$ number of sites and an operator string of length M, the storage requirements for the total simulation cell is $N_{s}(1+M)$ memory units, whereas for the spin-opstring data it is just $(N_{s}+M)$. As an example for a $12\times 12$ square lattice ($N_{s}=144$) with a typical $M=25000$, the spin-opstring requires about $140$ times less memory than the whole simulation cell. In terms of actual numbers, for $30000$ training samples of the two-dimensional (2D) $t_{2}-V_{1}$ model (discussed later in Sec. IV.1) on $12\times 12$ lattices, the total size of the spin-opstring data is $\sim 5$ Gigabytes whereas for full simulation cells, it is around $700$ Gigabytes. Therefore, with increasing system sizes, the memory requirements for the total simulation cell data render them practically unusable, making the spin-opstring data a clear choice.
@@ III Method
Before feeding the spin-opstring data to the machine, we preprocess the data by applying the following three steps. Firstly, the initial state $|\alpha\rangle$ is mapped from the spin basis to the occupation basis, such that $\uparrow$ spin at site $i$ ($\sigma(i)=1$) maps to the number operator $n_{i}=1$ and $\downarrow$ spin ($\sigma(i)=-1$) maps to $n_{i}=0$. Secondly, the operator string is normalized between values 0 and $1$, by dividing the operator string elements by the maximum possible value, which is $2N_{b}+1$, where $N_{b}$ is the total number of bonds in the system. Lastly, since the length of the operator string varies in SSE QMC simulations at different regions of the phase diagram, we determine the maximum length of the training as well as prediction data and apply a padding of zeroes at the end of the operator strings as required to equalize the length of the data. As a zero in the operator string signifies the identity operator, padding with zeroes at the end does not change the physical properties of the system encoded in the data.
@@
To analyze the 1D spin-opstring data, we have used the usual feed-forward neural network architecture with fully-connected (FC) layers and/or 1D convolutional neural network (CNN) layers. Each of these layers consists of an initial fully connected or convolution step, followed by batch normalization and a rectified linear unit (ReLU) activation function. For the final layer, a fully connected layer is used with a softmax activation function for classification tasks, and a ReLU activation for regression tasks, without batch normalization in both cases. During training, we have minimized the categorical cross-entropy loss function for classification tasks and the mean squared error for regression. The stochastic gradient descent learning process has been optimized by using the Adam algorithm [33]. The code has been implemented in Python by using the Keras library [34].
For the phase classification in the 2D $t_{2}-V_{1}$ model, we have used a fully-connected ML model with three layers: the first input layer consisting of a number of neurons equal to the spin-opstring data length, followed by a hidden layer of 32 neurons, and finally an output layer consisting of a few neurons, depending on the number of phases, with softmax activation. In contrast, for regression, the ML model consists of two fully-connected hidden layers, each with $64$ neurons and the final layer has a single neuron with a ReLU activation function.
Classification
model
Regression model
Input layer Conv1D $f_{n}=16$, $f_{s}=9$ BatchNorm ReLU Input layer
@@
For phase boundary detection in the 1D SSH model, we have used a CNN architecture. After the input layer, this architecture includes $4$ hidden CNN layers with $16,32,64$ and $128$ filters of sizes $9,5,5$ and $5$, respectively. Each CNN layer is followed by a MaxPool layer with a pool size of $2$ and a stride of $2$, which reduces the length of the 1D data by half. The output of the last CNN layer is then flattened and passed to a fully-connected layer with two neurons, corresponding to the two phases of the SSH model. The ML models used for the phase classification and regression tasks discussed above are schematically represented in Table 1.
For both of the domain adaptation tasks, we have employed 1D CNNs with a similar base architecture. Each network consists of three convolutional layers with kernel sizes of $9$, 55, and $5$, respectively. These layers are followed by batch normalization, ReLU activation, and a MaxPool layer with a pool size of $2$ and a stride of $2$. The model trained on $t_{2}-V_{1}$ uses $8$, $16$ and $32$ filters in the convolutional layers. The output from the last CNN is flattened and passed through a FC layer with $8$ neurons, followed by batch normalization, ReLU activation, which connects to a final FC layer of $2$ neurons with softmax activation. In contrast the model trained on the non-interacting SSH, employs $16$, $32$ and $64$ filters. Instead of flattening, after the final block, global average pooling is applied, followed by a FC layer consisting of $2$ neurons with softmax activation. During training in the SSH case, we employed dropout regularization after each Conv1D layer with a rate of $0.25$ to achieve better generalization.
For each of the ML models used in this study, we present our results obtained as an average of the predictions from 100 independently trained instances of the same model.
IV Applications : Phase classification and observable prediction
@@ IV.1 Two-dimensional $t_{2}-V_{1}$ model
We first apply our proposal to the 2D $t_{2}-V_{1}$ model which belongs to the family of extended Bose-Hubbard models in the hardcore limit. This simple model displays a rich phase diagram and is, in fact, the minimum model to realize lattice supersolidity of the checkerboard-type. Interestingly, the system also exhibits a line of mixed-order phase transitions with varying particle density which displays the extreme Thouless effect at half-filling. In the 2D $t_{2}-V_{1}$ model, a system of hard-core bosons (HCBs) is studied on a periodic square lattice described by the Hamiltonian,
@@
where $r_{i}=(x_{i},y_{i})$ represents the position of site $i$, $N_{s}$ the total number of sites, $\boldsymbol{Q}$ the wave vector and $\langle\cdots\rangle$ denotes ensemble average. For the calculation of the superfluid density using SSE method, the following formula, which involves the winding numbers $W_{x}$ and $W_{y}$ along $x$ and $y$-directions, is used,
@@
Figure 6: Comparison of ML models trained with spin-opstring and full simulation cell data for the SF-cS transition on $8\times 8$ lattices. The open symbols denote the probability for cS phase while the filled symbols stand for the SF phase. The errors on the points have been obtained from 20 independent training setups.
Although the spin-opstring data provides a memory-efficient compact representation of the full QMC simulation cell, it would be instructive at this stage to do a comparison of both of them. For this purpose, we consider the SF-cS transition on a $8\times 8$ lattice. We chose this smaller lattice because the size of the full simulation cell remains within feasible computational limits (each spin-opstring data has $64+5282=5346$ elements while each full simulation cell has $64\times(5282+1)=338112$ entries). Unlike in the case of spin-opstring, the use of a FC ML architecture is not viable for the full simulation cell as it leads to $\mathcal{O}(10^{7})$ trainable parameters even for this smaller lattice. We use instead a 2D CNN network leveraging the natural two-dimensional structure of the simulation cell where the horizontal direction is the site index of the spin-state and the vertical direction represents its time evolution.
The 2D CNN model has a similar architecture as described for the 1D SSH case in Table 1, having only three CNN layers with $8,16$ and $32$ filters respectively, followed by a FC classification layer. In Fig. 6, we show the predictions for the phase classification obtained from the two different ML models trained with the same number of data points. Both of them predict exactly the same transition point which perfectly matches with the QMC result. It appears that the 2D CNN model trained with the full simulation cell data is able to better reflect the sharp transition associated with the extreme Thouless effect. However, due to the larger size of the full simulation cell data, the training time per epoch on a Nvidia A100 GPU is about $6$ seconds whereas for the spin-opstring data, it is a mere $80$ milliseconds, a $75\times$ speedup.
@@
Figure 8: ML prediction of the SSH model performed on a 1D periodic chain with $32$ sites. As $v$ is varied from 0 to $1$, $w$ changes according to 1v1-v. The dashed line represents the phase boundary predicted by the ML model.